Suppose that all agents in the economy have the following utility function U(c,l)=( c(1-θ) /(1- θ ))-l  where c is consumption, l is the supply of labor, and θ a fixed parameter. Suppose that individuals only have labor income, with an hourly wage of w and a tax rate of t. Thus, the budget constraint of the agent is w(1-t) l=c . We will assume here that θ = 0.5 and w = 1.

The elasticity of the labor supply with respect to (1-t) is defined as e=(dl/d(1-t)) ((1-t)/l)  .

What is the value of this elasticity in our example?

Select one:

a.   None of the other answers is true.

b. 0.5

c. 0

d. 1.5

e. 1
Answer #1

to u C4,0)=CO) 129 320 80 MRS = au/al gulac co 123 oulac- (10) 00 (10) Yo = c-o at em MPS = wll-t) þcto = (1-4) as w=180 put 8 = 65 642= (1+) c= (1-62 then from Be. w(1-t) l = 6 1 C1-t) e = (1-t2 l= (+) 1.Ch y. dctt) (2) (10) optione.) (1-0) d

Answer option e 1